OpenStudy (anonymous):

the perimeter of a rectangle is 34yd , and the area of the rectangle is 72yd^2 . Find the dimensions of the rectangle.

OpenStudy (anonymous):

\[34=2l+2w\] and Solve the first equation for w (or l). \[34-2l=2w\] \[w=17-l\] Now substitute this result into the Area equation. \[72=(17-l)l\] \[72=17l-l^2\] \[l^2-17l+72=0\] Now factor that. \[(l-9)(l-8)=0\] Thus l = 8 or l= 9 . Either way the dimensions of the rectangle was 8 x 9. This would be the standard algebraic method of solving this equation. One could have solved it by guess and check with the hope that the answer was integer values.

OpenStudy (anonymous):

thank you